Given the aggregate consumption function C=0.9Y+100(where c is aggregate consumption and Y is aggregate income)

A. Find yhe marginal propensity to consume (mpc) and average propensity to consume (Apc)


1
Expert's answer
2021-02-03T20:25:37-0500

y=mx+c. than m=0.9. Slope of a curve gives us the value of mpc

MPS=0.9

APC=C/y. At y=10 C=109. then

APC=109/10=10.9